A man uses a rock as a fulcrum and lifts a 300-kg boulder with a metal bar 6.00 m long. Calculare the force necessary to lift the boulder if the fulcrum is:
(a) 1.00 m from where the bar contacts the rock.
(b) 3.00 m from where the bar contacts the rock.
1
Expert's answer
2018-01-29T10:19:07-0500
At the equilibrium the total torque around the pivot point (fulcrum) is zero, so
F_1 d_1=F_2 d_2
or
Fd_1=mgd_2
Thus
F=mg d_2/d_1
d_2=1 m, d_1=l-d_2=6-1=5 m
F=300×9.8×1/5=588 N
d_2=3 m, d_1=l-d_2=6-3=3 m
F=300×9.8×3/3=2940 N